OpenStudy (anonymous):

What is the x-intercept of the line with the equation -2x+1/2y=18

ganeshie8 (ganeshie8):

x-intercept is the point at which the line cuts the `x` axis

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

Notice that the y value is 0 when the line cuts the x axis

ganeshie8 (ganeshie8):

so you can set the y value to 0 and solve for x

ganeshie8 (ganeshie8):

-2x+1/2y=18 plugin y = 0 : -2x + 1/2(0) = 18 -2x = 18 x = ?

OpenStudy (anonymous):

so it's -9 ? ( sorry for late reply , my mums sick )